19. Gravetter/Wallnau/Forzano, Esse For a population with σ = 10, how large a sample is necessary to have a standard error that is: Less than 5 points? ntials Chapter 7 - End-of-chapter question 19 Aa Aa n > Less than 2 points? n > Less than 1 point? n >
Answer #1

Standard error = \frac{\sigma }{\sqrt{n}}

a) \frac{\sigma }{\sqrt{n}} < 5

\Rightarrow \frac{10 }{\sqrt{n}} < 5

\Rightarrow \sqrt{n} > \frac{10}{5}

\Rightarrow \sqrt{n} > 2

\Rightarrow n >4

b)

\frac{\sigma }{\sqrt{n}} < 2

\Rightarrow \frac{10 }{\sqrt{n}} < 2

\Rightarrow \sqrt{n} > \frac{10}{2}

\Rightarrow \sqrt{n} > 5

\Rightarrow n > 25

c)  \frac{\sigma }{\sqrt{n}} < 1

\Rightarrow \frac{10 }{\sqrt{n}} < 1

\Rightarrow \sqrt{n} > \frac{10}{1}

\Rightarrow \sqrt{n} > 10

\Rightarrow n > 100